OpenStudy (anonymous):

HELP PRECALC MEDALS 1.) Find an equation for the nth term of the arithmetic sequence. a14 = -33, a15 = 9 an = -579 + 42(n + 1) an = -579 + 42(n - 1) an = -579 - 42(n + 1) an = -579 - 42(n - 1) 2. Find an equation for the nth term of the arithmetic sequence. -15, -6, 3, 12, ... an = -15 + 9(n + 1) an = -15 x 9(n - 1) an = -15 + 9(n + 2) an = -15 + 9(n - 1) 3. Find an equation for the nth term of the sequence. -3, -12, -48, -192, an = 4 • -3n + 1 an = -3 • 4n - 1 an = -3 • 4n an = 4 • -3n

OpenStudy (misssunshinexxoxo):

Kindly separate these questions or tell us what you believe the answer is

OpenStudy (solomonzelman):

(Note: given two point. That can be an exponential function - geometric sequence, or a linear function - arithmetic sequence. However, I conclude from the answer choices that this #1 is arithm. sequence.)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle a_{14}=-33 }\), \(\large\color{black}{ \displaystyle a_{15}=9 }\) since we are talking about a geometric sequence \(\large\color{black}{ \displaystyle {\rm d}=a_{15}-a_{14}=9-(-33)=? }\)

OpenStudy (solomonzelman):

because in geometric sequence, d, the common difference between the terms is `(a term) - (a term before it)` so, \(\large\color{black}{ \displaystyle {\rm d}=a_{n-1}-a_{n}={\small(\rm in~this~case)}~~a_{15}-a_{14}=9-(-33)=? }\)

OpenStudy (anonymous):

@SolomonZelman an = -579 + 42(n - 1) (?)

OpenStudy (anonymous):

@SolomonZelman